OpenStudy (anonymous):

what is the lenght height and width of a rectanguar prism if they are in a ratio of 6:4:3 and the volume is 810

OpenStudy (amistre64):

what is 6*4*3?

OpenStudy (amistre64):

or rather; 6n * 4n * 3n = 810

OpenStudy (anonymous):

no its like the ratio, its 6:4:3

OpenStudy (anonymous):

so I need to find the measurments of each side by using those ratios

OpenStudy (amistre64):

the sides l:h:w have a ratio of 6:4:3 volume of a prism is defined as l*h*w = 810 the only difference between the actual lengths, and the ratio, are some constant scaled factor.

OpenStudy (amistre64):

therefore, we know: length = 6n height = 4n width = 3n and 6n*4n*3n = 810 solve for n

OpenStudy (anonymous):

ohh thank you but I made a mistake, the surface area is 810 and I need to find the volume of the rectangular prism

OpenStudy (amistre64):

...... same concepts, different "formula" tho can you recall a good formula for surface area?

OpenStudy (amistre64):

6n*4n*3n = Volume is still fine, and still have to determine the n value

OpenStudy (anonymous):

2ab+2bc+2ac right?

OpenStudy (amistre64):

sounds about right, but lets use lhw instead of abc

OpenStudy (anonymous):

so my equation would be 6n*4n*3n= 2lw+2wh+2hl ?

OpenStudy (amistre64):

2hl + 2hw + 2lw = 810 no, one formula is for surface area, and one is for volume. they will not generally equate to the same value

OpenStudy (amistre64):

lhw 643 2lw+2wh+2hl = 810 2(6n)(3n)+2(3n)(4n)+2(4n)(6n) = 810

OpenStudy (amistre64):

2(6n)(3n)+2(3n)(4n)+2(4n)(6n) = 810 36n^2 + 24n^2 + 48n^2 = 810 108n^2 = 810 n^2 = 810/108

OpenStudy (anonymous):

ohhhhh wait why would you square the numbers?

OpenStudy (anonymous):

so what does n represent?

OpenStudy (anonymous):

btw thank you so much for helping!!

OpenStudy (amistre64):

n represents the number that is scaling the ratio of 6:4:3 to its proper lengths

OpenStudy (anonymous):

so after I find n what do I do?

OpenStudy (amistre64):

i thought i covered that

OpenStudy (amistre64):

length is equal to 6n height is equal to 4n width is equal to 3n

OpenStudy (amistre64):

once you find n, you know that actual measurements and not just their ratio

OpenStudy (anonymous):

ooohhhh I get it wait so n would be 7.5?

OpenStudy (amistre64):

dunno, i aint got me calculator to find the sqrt of (810/108) but ill trust you on it

OpenStudy (anonymous):

wait why would you find the square root?

OpenStudy (amistre64):

you plug in the values for length width and height into the surface area formula and solve for n 2(6n)(3n)+2(3n)(4n)+2(4n)(6n) = 810 36n^2 + 24n^2 + 48n^2 = 810 108n^2 = 810 n^2 = 810/108

OpenStudy (anonymous):

wait is it possible to do a square root of a decimal?

OpenStudy (amistre64):

its possible to do the sqrt of any real number. it all depends on what you define as an acceptable output

OpenStudy (anonymous):

ohhh so I got 2.738612788 so would I just round it to the nearest tenth?

OpenStudy (amistre64):

rounding introduces a small error, the longer the decimal the smaller the error. so its up to you how far you want to take this in the end. my thought was to have 4 decies

OpenStudy (amistre64):

does the solution ask for an approxiamte, or an exact value?

OpenStudy (anonymous):

yeah it asks for an exact volume

OpenStudy (amistre64):

then do not round to anything :)

OpenStudy (amistre64):

you can simplify to:\[n=\sqrt{\frac{15}{2}}\]